OpenStudy (anonymous):

I need help on these problems.Simplify each expression if not possible write simplified. -6x+3x^2+10x^2

jimthompson5910 (jim_thompson5910):

3x^2 and 10x^2 are like terms combine them: 3x^2+10x^2 = 13x^2

jimthompson5910 (jim_thompson5910):

so the simplified expression is: -6x+13x^2 that can be written as 13x^2 - 6x

OpenStudy (anonymous):

Ty! I got tons of these problesm don't even know where to begin.

OpenStudy (anonymous):

Would you mind helping me out with some more? LOL. I'll just write em all in one message, or send you a picture how's that?

jimthompson5910 (jim_thompson5910):

I can help with a few more. Post them one at a time.

OpenStudy (anonymous):

Okay thanks. DO I need to make a new topic? Or sameo ne?

jimthompson5910 (jim_thompson5910):

either way is fine

OpenStudy (anonymous):

Kk. Next one is 2p+1/2q

jimthompson5910 (jim_thompson5910):

so \[\Large 2p + \frac{1}{2}q\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what are the instructions

OpenStudy (anonymous):

Same as before Simplify each expression if not possible write simplified.

jimthompson5910 (jim_thompson5910):

are there any like terms that you see?

jimthompson5910 (jim_thompson5910):

like terms are terms that have the same variable and exponents over the variable example: 2x and 3x are like terms (both have x) but 2x and 3y are not like terms (one has x, the other has y)

OpenStudy (anonymous):

Yes they are all on this problem, I think I understand what you mean.

jimthompson5910 (jim_thompson5910):

so in \[\Large 2p + \frac{1}{2}q\] we have no like terms because one has p, while the other has q. You cannot combine them

jimthompson5910 (jim_thompson5910):

you can't say for instance 2p + 3q = 5p or 2p+3q = 5q neither are correct or make sense

OpenStudy (anonymous):

Okay, see this is one confusing lol

OpenStudy (anonymous):

its simplied than i think lol

jimthompson5910 (jim_thompson5910):

keep practicing and you'll get the hang of it

OpenStudy (anonymous):

Wanna help with more? :p

OpenStudy (anonymous):

10xy-4(xy+xy)

jimthompson5910 (jim_thompson5910):

what are the like terms here?

OpenStudy (anonymous):

I got that one done, lol what is the abvolute value of 09? I dont get that.

jimthompson5910 (jim_thompson5910):

09 or -9 ?

OpenStudy (anonymous):

-9

jimthompson5910 (jim_thompson5910):

absolute value represents distance from that number to 0 on the number line how far is -9 from 0 on the number line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait oh duh lol would it be 9?

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

so |-9| = 9

jimthompson5910 (jim_thompson5910):

|-9| stands for "the absolute value of -9"

OpenStudy (anonymous):

I really appreciate you helping my btw

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

I onyl gotta catch up then my class is don lol

OpenStudy (anonymous):

Okay now this is changed.

OpenStudy (anonymous):

Let me see if i can find oyu the workshet just to show you one sec

OpenStudy (anonymous):

You know any of this?

jimthompson5910 (jim_thompson5910):

what is it asking exactly? some parts are cut off

OpenStudy (anonymous):

Hang on ill get other half lol

OpenStudy (anonymous):